OpenStudy (meilendurcer):

Lynne works at a bank and earns $9.75 per hour. If Lynne works 35 hours each week, which expression could be used to determine her total earnings for 1 year?

OpenStudy (anonymous):

nope...

OpenStudy (anonymous):

number of weeks per year * number of hours she works per week * earnings per hour

OpenStudy (anonymous):

First of all her want her earnings per year... Thus 1820 hrs in total (35 x 52) x $9.75. That's your equation.

OpenStudy (anonymous):

sorry I meant 'they want' not "her want".

OpenStudy (anonymous):

\[\frac{earnings}{year} = \frac{weeks}{year} * \frac{hours}{week} * \frac{earnings}{hour} = \frac{weeks*hours*earnings}{weeks*hours*year}\]